# Frequency Extraction from Invariant Flows

### v1

@@ Abstract
In non-degenerate integrable Hamiltonian systems, invariant tori can be parameterized equivalently by action variables or by their fundamental frequencies. We introduce an invariant-flow formulation for extracting fundamental frequencies of integrable Hamiltonian systems. By treating invariants as generators of commuting Hamiltonian flows, the frequencies are obtained from time-of-flight parameters along these flows, providing a direct alternative to action–angle constructions and spectral methods based on long time series. The approach yields an explicit numerical procedure that extends naturally to systems with multiple degrees of freedom. Its effectiveness is demonstrated using the McMillan map, where machine-precision accuracy is achieved.
@@
Integrable Hamiltonian systems possess as many independent, commuting invariants as degrees of freedom, and these invariants define smooth tori in the phase space. The quasi-periodic evolution on each torus is characterized by a set of fundamental frequencies. Extracting these frequencies directly from invariants provides an alternative to action–angle constructions, which are often difficult to obtain explicitly.
@@
The Liouville–Arnold theorem states that an integrable Hamiltonian system can be described in action–angle variables, in which the actions are constants of motion and the angles evolve as rigid rotations with fixed frequencies. In the one-degree-of-freedom (1DOF) case, this yields an explicit expression for the frequency (rotation number) in terms of the invariant. The Danilov equation expresses the frequency as [PhysRevAccelBeams.23.054001]
$\nu\left(K\right)=\left.{\int_{q_{0}}^{q_{1}}\left(\frac{\partial K}{\partial p}\right)^{-1}\mathrm{d}q}\right/{\oint\left(\frac{\partial K}{\partial p}\right)^{-1}\mathrm{d}q}$ (1)
Here $(q,p)$ are canonical coordinates, $(q_{0},p_{0})$ is an initial condition, and $(q_{1},p_{1})$ is its image under the one-turn map. The integrals are taken along the invariant curve $K(q,p)=\mathrm{const}$. However, extending this formulation beyond 1DOF is not straightforward.
Equation (1) can be interpreted as a time-of-flight ratio by viewing the invariant $K$ as a Hamiltonian generating a flow along its level set. Under this flow, the numerator gives the time required to transport the phase-space point from $(q_{0},p_{0})$ to its image $(q_{1},p_{1})$, while the denominator gives the period of the closed orbit on the invariant curve $K(q,p)=\mathrm{const}$. The frequency is the ratio of these two times.
@@
The Hamiltonian of a 2DOF integrable system can be written in action–angle variables $(\phi_{1},J_{1},\phi_{2},J_{2})$ as
$H=H(J_{1},J_{2}).$ (2)
@@
$\nu_{i}=\frac{1}{2\pi}\frac{\partial H}{\partial J_{i}},\qquad i=1,2.$ (3)
The explicit form of $H$ and the associated actions are generally unavailable. Instead, the dynamics is accessed through the discrete symplectic map generated by the Hamiltonian flow,
$\Phi_{H}^{t}=\exp\!\left(-t\colon H\colon\right),\qquad t=0,1,2,\ldots,$ (4)
where $\colon H\colon f=[H,f]$ denotes the Poisson bracket for an arbitrary function $f$.
@@
$K_{1}=K_{1}(J_{1},J_{2}),\qquad K_{2}=K_{2}(J_{1},J_{2}),$ (5)
@@
$H,f]=\frac{\partial H}{\partial K_{1}}[K_{1},f]+\frac{\partial H}{\partial K_{2}}[K_{2},f],$ (6)
@@
$\colon H\colon=\frac{\partial H}{\partial K_{1}}\colon K_{1}\colon+\frac{\partial H}{\partial K_{2}}\colon K_{2}\colon$. (7)
Since the flows generated by $K_{1}$ and $K_{2}$ commute, the Baker–Campbell–Hausdorff (BCH) formula gives the factorization of the Hamiltonian flow,
$\exp\left(-t\colon H\colon\right)=\exp$ $\!\left[-t\!\left(\frac{\partial H}{\partial K_{1}}\right)\!\colon K_{1}\colon\right$] (8)
$\exp\!\left[-t\!\left(\frac{\partial H}{\partial K_{2}}\right)\!\colon K_{2}\colon\right$].
@@
$\tau_{1}(t)=t\left(\frac{\partial H}{\partial K_{1}}\right),\qquad\tau_{2}(t)=t\left(\frac{\partial H}{\partial K_{2}}\right),$ (9)
the $H$-flow can be written compactly as a composition of invariant flows,
$\Phi_{H}^{t}=\Phi_{K_{1}}^{\tau_{1}}\Phi_{K_{2}}^{\tau_{2}}=\Phi_{K_{2}}^{\tau_{2}}\Phi_{K_{1}}^{\tau_{1}}.$ (10)
The flows generated by $H$, $K_{1}$, and $K_{2}$ preserve the actions and induce rigid rotations but with distinct angular velocities. For the factorized representation to reproduce the same one-turn evolution ($t=1$ in Eq. 4), the angle increments generated by the composition of the $K_{1}$ and $K_{2}$ flows must match those produced by the $H$-flow. Consistency therefore requires
$\begin{pmatrix}\partial K_{1}/\partial J_{1}&\partial K_{2}/\partial J_{1}\\ \partial K_{1}/\partial J_{2}&\partial K_{2}/\partial J_{2}\end{pmatrix}\begin{pmatrix}\tau_{1}\\ \tau_{2}\end{pmatrix}=\begin{pmatrix}\partial H/\partial J_{1}\\ \partial H/\partial J_{2}\end{pmatrix}$. (11)
The motion is confined to a smooth two-dimensional invariant torus. In the non-resonant case, the trajectory on a given torus is quasi-periodic and densely fills the torus. The flows generated by $H$, $K_{1}$, and $K_{2}$ remain on the same torus and translate the phase-space point along directions defined by their Hamiltonian vector fields $(\dot{x},\dot{p}_{x},\dot{y},\dot{p}_{y})$. Because $K_{1}$ and $K_{2}$ are independent invariants that Poisson–commute, their flows generate two linearly independent directions on the torus. By composing these flows with appropriate time-of-flight parameters, any point on the same invariant torus can be reached. In particular, suitable combinations generate closed loops on the torus,
$\Phi_{K_{1}}^{T_{1}}\Phi_{K_{2}}^{T_{2}}=\exp\!\left[-2\pi:\left(m_{1}J_{1}+m_{2}J_{2}\right):\right],$ (12)
where $m_{1,2}$ are integers specifying the winding numbers associated with the actions $J_{1,2}$, and $T_{1,2}$ are the corresponding time-of-flight parameters under the $K_{1,2}$ flows. Consistency with Eq. (11) then yields
$\begin{pmatrix}\partial K_{1}/\partial J_{1}&\partial K_{2}/\partial J_{1}\\ \partial K_{1}/\partial J_{2}&\partial K_{2}/\partial J_{2}\end{pmatrix}\begin{pmatrix}T_{1}\\ T_{2}\end{pmatrix}=2\pi\begin{pmatrix}m_{1}\\ m_{2}\end{pmatrix}$. (13)
@@
$\begin{pmatrix}\partial K_{1}/\partial J_{1}&\partial K_{2}/\partial J_{1}\\ \partial K_{1}/\partial J_{2}&\partial K_{2}/\partial J_{2}\end{pmatrix}=2\pi\begin{pmatrix}m_{11}&m_{12}\\ m_{21}&m_{22}\end{pmatrix}\begin{pmatrix}T_{11}&T_{12}\\ T_{21}&T_{22}\end{pmatrix}^{-1}.$ (14)
@@
The parameters $\tau_{1}$ and $\tau_{2}$ are determined from symplectic integration of the invariant flows. Let $\mathbf{z}=(x,p_{x},y,p_{y})^{\mathrm{T}}$ denote the phase-space vector and $\mathbf{z}_{0}$ an initial point. Applying the one-turn map yields the image $\Phi_{H}(\mathbf{z}_{0})$. For trial values of $\tau_{1}$ and $\tau_{2}$, a reconstructed point is obtained by composing the invariant flows,
$\mathbf{z}\left(\tau_{1},\tau_{2}\right)=\Phi_{K_{2}}^{\tau_{2}}\left(\Phi_{K_{1}}^{\tau_{1}}\left(\mathbf{z}_{0}\right)\right)=\Phi_{K_{1}}^{\tau_{1}}\left(\Phi_{K_{2}}^{\tau_{2}}\left(\mathbf{z}_{0}\right)\right)$ (16)
@@
$\mathcal{L}(\tau_{1},\tau_{2})=\bigl\|\mathbf{z}\left(\tau_{1},\tau_{2}\right)-\Phi_{H}(\mathbf{z}_{0})\bigr\|.$ (17)
@@
$\frac{\partial\mathbf{z}}{\partial\tau_{i}}=[\mathbf{z},K_{i}]\bigg|_{\mathbf{z}=\mathbf{z}(\tau_{1},\tau_{2})},\qquad i=1,2.$ (18)
@@
$\mathcal{L}(T_{1},T_{2})=\bigl\|\mathbf{z}\left(T_{1},T_{2}\right)-\mathbf{z}_{0}\bigr\|.$ (19)
@@
Even restricting the frequencies to the physically relevant range $0<\nu_{1,2}<0.5$ does not uniquely determine the integer matrix $\mathbf{m}$. To fix this ambiguity, a coarse estimate can be obtained analytically from the linear one-turn map. Although such estimates are not sufficiently precise for high-accuracy frequency determination, they are adequate to discriminate among the finite set of candidate integer matrices. Only one choice yields a frequency vector consistent with this estimate, thereby fixing the cycle basis unambiguously.
@@
$\begin{gathered}K_{1}=x^{2}+y^{2}+p_{x}^{2}+p_{y}^{2}-a(xp_{x}+yp_{y})+b(xp_{x}+yp_{y})^{2},\\ K_{2}=xp_{y}-yp_{x}.\end{gathered}$ (25)
@@
An additional advantage of the invariant-flow formulation is that it does not require analytic expressions for the invariants themselves. The method relies only on the evaluation of the Hamiltonian vector fields generated by the invariants, $\mathbf{z},K_{i}]$, which depend on derivatives of the invariants with respect to phase-space coordinates. This makes the approach naturally compatible with data-driven or neural-network-based representations of invariants, where gradients are readily available through automatic differentiation.
In contrast to approaches that attempt to learn action–angle variables directly, the present framework does not require enforcing global angle periodicity or canonical structure during training. Once approximate invariants are available, the frequencies can be extracted directly from the corresponding invariant flows. This suggests a natural route for combining modern data-driven techniques with geometric frequency analysis.

### v2

@@ Abstract
In non-degenerate integrable Hamiltonian systems, invariant tori can be parameterized equivalently by action variables or by their fundamental frequencies. We introduce an invariant-flow formulation for extracting fundamental frequencies of integrable Hamiltonian systems. By treating invariants as generators of commuting Hamiltonian flows, the frequencies are obtained from time-of-flight parameters along these flows, providing a direct alternative to action-angle constructions and spectral methods based on long time series. The approach yields an explicit numerical procedure that extends naturally to systems with multiple degrees of freedom. Its effectiveness is demonstrated using the McMillan map, where machine-precision accuracy is achieved.
@@
Integrable Hamiltonian systems possess as many independent, commuting invariants as degrees of freedom, and these invariants define smooth tori in the phase space. The quasi-periodic evolution on each torus is characterized by a set of fundamental frequencies. Extracting these frequencies directly from invariants provides an alternative to action-angle constructions, which are often difficult to obtain explicitly.
@@
The Liouville-Arnold theorem states that an integrable Hamiltonian system can be described in action-angle variables, in which the actions are constants of motion and the angles evolve as rigid rotations with fixed frequencies. In the one-degree-of-freedom (1DOF) case, this yields an explicit expression for the frequency (rotation number) in terms of the invariant. The Danilov equation expresses the frequency as [PhysRevAccelBeams.23.054001]
$\nu\left(\mathcal{K}\right)=\left.{\int_{q_{0}}^{q_{1}}\left(\frac{\partial\mathcal{K}}{\partial p}\right)^{-1}\mathrm{d}q}\right/{\oint\left(\frac{\partial\mathcal{K}}{\partial p}\right)^{-1}\mathrm{d}q}$ (1)
Here $(q,p)$ are canonical coordinates, $(q_{0},p_{0})$ is an initial condition, and $(q_{1},p_{1})$ is its image under the one-turn map. The integrals are taken along the invariant curve $\mathcal{K}(q,p)=\mathrm{const}$. However, extending this formulation beyond 1DOF is not straightforward.
Equation (1) can be interpreted as a time-of-flight ratio by viewing the invariant $\mathcal{K}$ as a Hamiltonian generating a flow along its level set. Under this flow, the numerator gives the time required to transport the phase-space point from $(q_{0},p_{0})$ to its image $(q_{1},p_{1})$, while the denominator gives the period of the closed orbit on the invariant curve $\mathcal{K}(q,p)=\mathrm{const}$. The frequency is the ratio of these two times.
@@
The Hamiltonian of a 2DOF integrable system can be written in action-angle variables $(\phi_{1},J_{1},\phi_{2},J_{2})$ as
$\mathcal{H}=\mathcal{H}(J_{1},J_{2}).$ (2)
@@
$\nu_{i}=\frac{1}{2\pi}\frac{\partial\mathcal{H}}{\partial J_{i}},\qquad i=1,2.$ (3)
The explicit form of $\mathcal{H}$ and the associated actions are generally unavailable. Instead, the dynamics is accessed through the discrete symplectic map generated by the Hamiltonian flow,
$\Phi_{\mathcal{H}}^{t}=\exp\!\left(-t\colon\mathcal{H}\colon\right),\qquad t=0,1,2,\ldots,$ (4)
where $\colon\mathcal{H}\colon f=[\mathcal{H},f]$ denotes the Poisson bracket for an arbitrary function $f$.
@@
$\mathcal{K}_{1}=\mathcal{K}_{1}(J_{1},J_{2}),\qquad\mathcal{K}_{2}=\mathcal{K}_{2}(J_{1},J_{2}),$ (5)
@@
$\mathcal{H},f]=\frac{\partial\mathcal{H}}{\partial\mathcal{K}_{1}}[\mathcal{K}_{1},f]+\frac{\partial\mathcal{H}}{\partial\mathcal{K}_{2}}[\mathcal{K}_{2},f],$ (6)
@@
$\colon\mathcal{H}\colon=\frac{\partial\mathcal{H}}{\partial\mathcal{K}_{1}}\colon\mathcal{K}_{1}\colon+\frac{\partial\mathcal{H}}{\partial\mathcal{K}_{2}}\colon\mathcal{K}_{2}\colon$. (7)
Since the flows generated by $\mathcal{K}_{1}$ and $\mathcal{K}_{2}$ commute, the Baker–Campbell–Hausdorff (BCH) formula gives the factorization of the Hamiltonian flow,
$\exp\left(-t\colon\mathcal{H}\colon\right)=\exp$ $\!\left[-t\!\left(\frac{\partial\mathcal{H}}{\partial\mathcal{K}_{1}}\right)\!\colon\mathcal{K}_{1}\colon\right$] (8)
$\exp\!\left[-t\!\left(\frac{\partial\mathcal{H}}{\partial\mathcal{K}_{2}}\right)\!\colon\mathcal{K}_{2}\colon\right$].
@@
$\tau_{1}(t)=t\left(\frac{\partial\mathcal{H}}{\partial\mathcal{K}_{1}}\right),\qquad\tau_{2}(t)=t\left(\frac{\partial\mathcal{H}}{\partial\mathcal{K}_{2}}\right),$ (9)
the $\mathcal{H}$-flow can be written compactly as a composition of invariant flows,
$\Phi_{\mathcal{H}}^{t}=\Phi_{\mathcal{K}_{1}}^{\tau_{1}}\Phi_{\mathcal{K}_{2}}^{\tau_{2}}=\Phi_{\mathcal{K}_{2}}^{\tau_{2}}\Phi_{\mathcal{K}_{1}}^{\tau_{1}}.$ (10)
The flows generated by $\mathcal{H}$, $\mathcal{K}_{1}$, and $\mathcal{K}_{2}$ preserve the actions and induce rigid rotations but with distinct angular velocities. For the factorized representation to reproduce the same one-turn evolution ($t=1$ in Eq. 4), the angle increments generated by the composition of the $\mathcal{K}_{1}$ and $\mathcal{K}_{2}$ flows must match those produced by the $\mathcal{H}$-flow. Consistency therefore requires
$\begin{pmatrix}\partial\mathcal{K}_{1}/\partial J_{1}&\partial\mathcal{K}_{2}/\partial J_{1}\\ \partial\mathcal{K}_{1}/\partial J_{2}&\partial\mathcal{K}_{2}/\partial J_{2}\end{pmatrix}\begin{pmatrix}\tau_{1}\\ \tau_{2}\end{pmatrix}=\begin{pmatrix}\partial\mathcal{H}/\partial J_{1}\\ \partial\mathcal{H}/\partial J_{2}\end{pmatrix}$. (11)
The motion is confined to a smooth two-dimensional invariant torus. In the non-resonant case, the trajectory on a given torus is quasi-periodic and densely fills the torus. The flows generated by $\mathcal{H}$, $\mathcal{K}_{1}$, and $\mathcal{K}_{2}$ remain on the same torus and translate the phase-space point along directions defined by their Hamiltonian vector fields $(\dot{x},\dot{p}_{x},\dot{y},\dot{p}_{y})$. Because $\mathcal{K}_{1}$ and $\mathcal{K}_{2}$ are independent invariants that Poisson-commute, their flows generate two linearly independent directions on the torus. By composing these flows with appropriate time-of-flight parameters, any point on the same invariant torus can be reached. In particular, suitable combinations generate closed loops on the torus,
$\Phi_{\mathcal{K}_{1}}^{T_{1}}\Phi_{\mathcal{K}_{2}}^{T_{2}}=\exp\!\left[-2\pi:\left(m_{1}J_{1}+m_{2}J_{2}\right):\right],$ (12)
where $m_{1,2}$ are integers specifying the winding numbers associated with the actions $J_{1,2}$, and $T_{1,2}$ are the corresponding time-of-flight parameters under the $\mathcal{K}_{1,2}$ flows. Consistency with Eq. (11) then yields
$\begin{pmatrix}\partial\mathcal{K}_{1}/\partial J_{1}&\partial\mathcal{K}_{2}/\partial J_{1}\\ \partial\mathcal{K}_{1}/\partial J_{2}&\partial\mathcal{K}_{2}/\partial J_{2}\end{pmatrix}\begin{pmatrix}T_{1}\\ T_{2}\end{pmatrix}=2\pi\begin{pmatrix}m_{1}\\ m_{2}\end{pmatrix}$. (13)
@@
$\begin{pmatrix}\partial\mathcal{K}_{1}/\partial J_{1}&\partial\mathcal{K}_{2}/\partial J_{1}\\ \partial\mathcal{K}_{1}/\partial J_{2}&\partial\mathcal{K}_{2}/\partial J_{2}\end{pmatrix}=2\pi\begin{pmatrix}m_{11}&m_{12}\\ m_{21}&m_{22}\end{pmatrix}\begin{pmatrix}T_{11}&T_{12}\\ T_{21}&T_{22}\end{pmatrix}^{-1}.$ (14)
@@
The parameters $\tau_{1}$ and $\tau_{2}$ are determined from symplectic integration of the invariant flows. Let $\mathbf{z}=(x,p_{x},y,p_{y})^{\mathrm{T}}$ denote the phase-space vector and $\mathbf{z}_{0}$ an initial point. Applying the one-turn map yields the image $\Phi_{\mathcal{H}}(\mathbf{z}_{0})$. For trial values of $\tau_{1}$ and $\tau_{2}$, a reconstructed point is obtained by composing the invariant flows,
$\mathbf{z}\left(\tau_{1},\tau_{2}\right)=\Phi_{\mathcal{K}_{2}}^{\tau_{2}}\left(\Phi_{\mathcal{K}_{1}}^{\tau_{1}}\left(\mathbf{z}_{0}\right)\right)=\Phi_{\mathcal{K}_{1}}^{\tau_{1}}\left(\Phi_{\mathcal{K}_{2}}^{\tau_{2}}\left(\mathbf{z}_{0}\right)\right)$ (16)
@@
${L}(\tau_{1},\tau_{2})=\bigl\|\mathbf{z}\left(\tau_{1},\tau_{2}\right)-\Phi_{\mathcal{H}}(\mathbf{z}_{0})\bigr\|.$ (17)
@@
$\frac{\partial\mathbf{z}}{\partial\tau_{i}}=[\mathbf{z},\mathcal{K}_{i}]\bigg|_{\mathbf{z}=\mathbf{z}(\tau_{1},\tau_{2})},\qquad i=1,2.$ (18)
@@
${L}(T_{1},T_{2})=\bigl\|\mathbf{z}\left(T_{1},T_{2}\right)-\mathbf{z}_{0}\bigr\|.$ (19)
@@
Even restricting the frequencies to the range $0<\nu_{1,2}<0.5$ does not uniquely determine the integer matrix $\mathbf{m}$. To fix this ambiguity, a coarse estimate can be obtained analytically from the linear one-turn map. Although such estimates are not sufficiently precise for high-accuracy frequency determination, they are adequate to discriminate among the finite set of candidate integer matrices. Only one choice yields a frequency vector consistent with this estimate, thereby fixing the cycle basis unambiguously.
@@
$\begin{gathered}\mathcal{K}_{1}=x^{2}+y^{2}+p_{x}^{2}+p_{y}^{2}-a(xp_{x}+yp_{y})+b(xp_{x}+yp_{y})^{2},\\ \mathcal{K}_{2}=xp_{y}-yp_{x}.\end{gathered}$ (25)
@@
An additional advantage of the invariant-flow formulation is that it does not require analytic expressions for the invariants themselves. The method relies only on the evaluation of the Hamiltonian vector fields generated by the invariants, $\mathbf{z},\mathcal{K}_{i}]$, which depend on derivatives of the invariants with respect to phase-space coordinates. This makes the approach naturally compatible with data-driven or neural-network-based representations of invariants, where gradients are readily available through automatic differentiation.
In contrast to approaches that attempt to learn action-angle variables directly, the present framework does not require enforcing global angle periodicity or canonical structure during training. Once approximate invariants are available, the frequencies can be extracted directly from the corresponding invariant flows. This suggests a natural route for combining modern data-driven techniques with geometric frequency analysis.